OpenStudy (anonymous):

How to factor y^4-16

OpenStudy (anonymous):

use A^2 - B^2 = (A+B)(A-B)

OpenStudy (whpalmer4):

And don't stop there! After the first factoring using the "difference of squares", one of the factors will also be a difference of squares and can be factored in the same fashion.

OpenStudy (anonymous):

y^4 is a square of a square and so is 16. any 4th power is a square of a square so wouldn't it be that both factors are squares after you factor the difference because the expression y^4-16 is equivalent to y^4-2^4?

OpenStudy (whpalmer4):

\[y^4 - 16 = (y^2)^2 - (4)^2 = (y^2-4)(y^2+4)\]But \(y^2-4\) is also a difference of squares: \[(y^2-4) = (y)^2-(2)^2 = (y-2)(y+2)\] \(y^2+4\) is NOT a difference of squares, so we have no easy way to factor it, unlike for \(y^2-4\). So our final factoring is \[y^4-16 = (y^2-4)(y^2+4) = (y-2)(y+2)(y^2+4)\]

OpenStudy (anonymous):

yeah it is a sum of squares but you can still factor that. y^2 + 4 = (y+2i)(y-2i)

OpenStudy (whpalmer4):

Only if you assume that the student is factoring over complex numbers and not just reals. In my experience, that isn't true; students getting factoring problems like this are not dealing with complex numbers.

OpenStudy (anonymous):

but it is good to learn that factoring a sum of squares gives you something similar to the difference of squares but with complex numbers. It is also good to learn that not only is i = sqrt(-1) but that -i = -sqrt(-1) as well